OpenStudy (anonymous):

hello-- how do i do (3^-2)^(5/2) please i want to learn how to do it thanks God bless you all

OpenStudy (anonymous):

\[(3^{-2})^{5/2}\]?

jhonyy9 (jhonyy9):

3^-2 = 1/9 (1/9)^5/2 = radical(1/9)^5 =(1/9)^2radical(1/9)=1/81 (radical(1/9) = =(1/81)(1/3)=1/3^5

jhonyy9 (jhonyy9):

=3^(-5)